General Physics

square root of-

5^2 + 5^2 + 2*5*5*cos60 (5^2 means five square)

= 2*5*cos30

how did this come, espacially 2 and cos30???????

Sameer Ahmed
15 Years agoGrade

3 Answers

Sanjeev Malik IIT BHU
15 Years ago

52+52+2*52*1/2  (cos60=1/2)  = 3*52

its square root =5*square root(3)

now,

2*5*cos30=2*5*root(3)/2   (cos30=root(3)/2)  = 5*root(3)

hence both are equal

athresh karanam
15 Years ago

cos60=1/2.so the given expression becomes

sq. root of 52+52+5*5=5*31/2=2*5*31/2=2*5*cos30..........(because cos30=31/2/2)

 

Aman Ahmad
5 Years ago
I wanted to know the answer too so I researched  I know this is from Concept Of Physics 1 Chapter 2, So If we modify the formula sqrt A^2 +B^2+2AB Cos’theta’ for A=B we get 2Acos’theta’/2 and if we put the values in it we get 2x5cos30’degree’ where A=B.
I hope you get the answer.